# Distributional Preference Learning: Understanding and Accounting for Hidden Context in RLHF

### v1

@@ 1 Introduction
Beyond identifying potential instances of relevant hidden context, our experiments indicate that DPL can be used to develop guardrails that protect against jailbreaks. (Wei et al., 2023) showed that many jailbreaks succeed by pitting the helpfulness and harmlessness objectives of chatbots against one another. This means that some jailbreaks can be understood as a consequence of hidden context. As a result, it is possible to detect this class of jailbreaks by leveraging the distribution of utilities we get from DPL. In particular, risk-aversion with respect to the distribution of learned utilities can dramatically reduce the rate at which the preference model prefers jailbroken responses. This is because DPL models capture the disagreement in the training dataset between the two objectives. Thus, risk-aversion penalizes responses that cause the two objectives to diverge.
@@ 2 Setting and Related Work
In standard preference learning, the goal is to estimate an unknown utility function $u:\mathcal{A}\to\mathbb{R}$ which assigns a utility value to each of a finite set of alternatives $\mathcal{A}$. For instance, in the case of a chatbot, the alternatives could be the possible responses to a prompt, and the utility function would describe how much a particular response is preferred. To estimate $u$, we observe the outcome of comparisons between pairs of alternatives $(a,b)$. We assume there is a fixed probability for any pair of alternatives $(a,b)$ that $a$ will be preferred to $b$; we denote this probability $p_{u}(a,b)$ and assume that $p_{u}(a,b)+p_{u}(b,a)=1$; that is, the order in which the alternatives are presented does not matter. In the ideal case, comparison outcomes would exactly reflect the utility function, i.e., $p_{u}(a,b)=\textbf{1}\{u(a)>u(b)\}$. Realistically, however, preference comparison data never exactly follows a single utility function. To account for the fact that people are noisy and/or inconsistent in their feedback, a common assumption is that instead preference comparisons are made according to a Bradley-Terry-Luce (BTL) model (Rajkumar & Agarwal, 2014), also sometimes known as Boltzmann-rational model (Jeon et al., 2020), model:
@@
The most commonly used method for estimating the utility function $u$ from preference data is to fit the maximum likelihood estimator (MLE) under the BTL model in (1). To derive the MLE, we consider the limit of infinite data and assume that preference comparisons are elicited for uniformly randomly selected pairs of alternatives. Under these conditions, the regularized MLE for the utility function $\hat{u}$ is given by minimizing the following loss function:
@@ 2.1 Hidden Context
$p_{u,\mathcal{D}_{z}}(a,b)=\mathbb{E}_{z\sim\mathcal{D}_{z}}\left[{O_{u}}(a,b,z)\right]\ ,\ {O_{u}}(a,b,z)=\begin{cases}1/2\quad\text{if }u(a,z)=u(b,z)\\ \mathbf{1}\{u(a,z)>u(b,z)\}\quad\text{o.w.}\end{cases}$ (4)
@@ 2.2 Related Work
The challenges of preference learning that we group as cases of “hidden context” have also been studied individually. There has been some work on explicitly modeling annotator disagreement (Fleisig et al., 2023; Baumler et al., 2023); these methods could be considered related to our proposed distributional preference learning (DPL) approach since they also aim to characterize alternatives where there is disagreement in the data. However, Fleisig et al. (2023) additionally require annotator information to be used for training. Baumler et al. (2023) consider learning from scalar (e.g., Likert scale) feedback rather than preference comparisons. Thus, DPL requires simpler data than both approaches and can also be applied to other types of hidden context.
Some work has studied the effects of human irrationality or non-BTL models of human behavior on preference learning (Bobu et al., 2020; Lee et al., 2021; Laidlaw & Russell, 2021; Knox et al., 2022; Laidlaw & Dragan, 2022). As we argue above, these deviations from the BTL model can be modeled in the hidden context setting by considering noise variables or human model parameters as hidden context. Zhuang & Hadfield-Menell (2020) consider the consequences of incorrectly optimizing a combination of objectives; this is related to Section 5, where we study LLM chatbots optimized to be both helpful and harmless.
Related to our connections with social choice theory in Section 3, some previous work has associated preference or reward learning with concepts in economics, such as voting rules (Conitzer & Sandholm, 2005), incentive compatibility (Echenique & Prasad, 2019), and mechanism design (Fickinger et al., 2020). RLHF in particular has been connected with literature on social choice, primarily through the lens of impossibility results (Mishra, 2023). To this end, there has been growing interest in designing systems that incorporate appropriate social objectives, or account for multiple stakeholders (Lambert et al., 2023), while eliciting preference in the context of recommender systems (Jia et al., 2023) and language modelling (Fish et al., 2023).
Similarly to our experiments in Section 5, Dai et al. (2023) study the conflicting helpfulness and harmlessness objectives in RLHF. They propose a method, safe RLHF, for explicitly modeling these objectives separately; in contrast, our approach of combining risk-averse optimization with distributional preference learning (DPL) does not require explicit tracking of different objectives. Furthermore, unlike safe RLHF, our approach is also applicable to the other settings we group under “hidden context,” like learning from a population with diverse preferences.
@@ 3 Theoretical Analysis
[…]
@@ 3.1 Comparison to expected utility and least-squares regression
The fact that least-squares utility regression converges to $\hat{u}=\bar{u}$ shows that, in some sense, it gracefully degrades in the presence of hidden context. Although there are problems with optimizing expected utility, it is at least a well-understood method of aggregating utilities over the hidden contexts and it has desirable decision-theoretic properties. When does preference learning with hidden context converge to the expected utility as well? We now formally characterize some situations where preference learning behaves similarly to utility regression and also describe cases when they behave completely differently.
@@
Many noise distributions, such as uniform and normal distributions, clearly satisfy the assumptions of Theorem 3.2. Thus, as long as the noise caused by hidden context does not vary across alternatives and is not too unusual, we generally expect that preference learning will give a utility function equivalent to the expected utility—similarly to least-squares regression.
@@
[…]
@@ 3.2 Connections to social choice theory
When training on comparison data coming from many agents, each with their own preferences, preference learning aggregates all their feedback into a single utility function. As we described in Section 2, this is a case where the identity of the annotator is a missing variable: it affects the comparison outcomes, but is unseen by the preference learning algorithm. Social choice theory studies methods for aggregating preferences from a population. Thus, it can provide a lens through which to understand this particular case of preference learning with hidden contexts.
@@ 4 Distributional Preference Learning
Our alternative preference learning methods, which we call distributional preference learning (DPL), output a distribution over possible utilities for each alternative rather than a single value (Figure 3). In particular, we learn a mapping $\smash{\hat{\mathcal{D}}}:\mathcal{A}\to\Delta(\mathbb{R})$ from alternatives to distributions over utilities to estimate the distribution of $u(a,z)$ when $z\sim\mathcal{D}_{z}$. We consider two variants, each of which parameterizes the distribution $\mathcal{D}_{z}(a)$ in a different way.
First, the mean-and-variance model learns two functions $\hat{\mu}:\mathcal{A}\to\mathbb{R}$ and $\hat{\sigma}:\mathcal{A}\to[0,\infty)$, parameterizing the distribution over utilities as $\smash{\hat{\mathcal{D}}}(a)=\mathcal{N}\left(\hat{\mu}(a),\hat{\sigma}(a)^{2}\right)$. Second, in the categorical model, we choose $n$ evenly spaced utility values $u_{1}<u_{2}<\ldots<u_{n}$, and then parameterize the distribution as the probabilities of each of those utilities $\hat{p}(u_{i}\mid a)$ for $i\in\{1,\dots,n\}$. We train the distributional preference models by maximizing the likelihood of the data given the model $p_{\smash{\hat{\mathcal{D}}}}(a,b)=\mathbb{E}\left[{O}(u_{a},u_{b})|u_{a}\sim\smash{\hat{\mathcal{D}}}(a),u_{b}\sim\smash{\hat{\mathcal{D}}}(b)\right]$. Concretely, for the mean-and-variance model, the loss for a single preference comparison where alternative $a$ is preferred to $b$ is the negative log probability that $u_{a}-u_{b}>0:$
@@ 5 Case Study: Competing Objectives in RLHF
In this section, we evaluate DPL’s ability to identify hidden context through a case study on large language model (LLM)-based reward models. Chatbots like GPT-4 and Claude are trained by learning a human reward model and then optimizing it via reinforcement learning, together referred to as RLHF. There are lots of ways that hidden context can affect reward models. In order to evaluate the ability of DPL methods to identify hidden context, we use the HH-RLHF dataset (Bai et al., 2022a). For this dataset, raters were separately asked to provide preferences on whether responses were helpful or harmful. This allows us to determine if DPL can automatically detect this context.
@@ B.2 BTL and Borda Count
[…]

### v2

@@ 1 Introduction
Beyond identifying potential instances of relevant hidden context, our experiments indicate that DPL can be used to develop guardrails that protect against jailbreaks. Wei et al. (2023) showed that many jailbreaks succeed by pitting the helpfulness and harmlessness objectives of chatbots against one another. This means that some jailbreaks can be understood as a consequence of hidden context. As a result, it is possible to detect this class of jailbreaks by leveraging the distribution of utilities we get from DPL. In particular, risk-aversion with respect to the distribution of learned utilities can dramatically reduce the rate at which the preference model prefers jailbroken responses. This is because DPL models capture the disagreement in the training dataset between the two objectives. Thus, risk-aversion penalizes responses that cause the two objectives to diverge.
@@ 2 Setting and Related Work
In standard preference learning, the goal is to estimate an unknown utility function $u:\mathcal{A}\to\mathbb{R}$ which assigns a utility value to each of a finite set of alternatives $\mathcal{A}$. For instance, in the case of a chatbot, the alternatives could be the possible responses to a prompt, and the utility function would describe how much a particular response is preferred. To estimate $u$, we observe the outcome of comparisons between pairs of alternatives $(a,b)$. We assume there is a fixed probability for any pair of alternatives $(a,b)$ that $a$ will be preferred to $b$; we denote this probability $p_{u}(a,b)$ and assume that $p_{u}(a,b)+p_{u}(b,a)=1$; that is, the order in which the alternatives are presented does not matter. In the ideal case, comparison outcomes would exactly reflect the utility function, i.e., $p_{u}(a,b)=\textbf{1}\{u(a)>u(b)\}$. Realistically, however, preference comparison data never exactly follows a single utility function. To account for the fact that people are noisy and/or inconsistent in their feedback, a common assumption is that instead preference comparisons are made according to a Bradley-Terry-Luce (BTL) model (Rajkumar & Agarwal, 2014), also sometimes known as Boltzmann-rational model (Jeon et al., 2020):
@@
The most commonly used method for estimating the utility function $u$ from preference data is to fit the maximum likelihood estimator (MLE) under the BTL model in (1). To derive the MLE, we consider the limit of infinite data and assume that preference comparisons are elicited for uniformly randomly selected pairs of alternatives. The MLE for the utility function $\hat{u}$ is given by minimizing the following loss function:
@@ 2.1 Hidden Context
$p_{u,\mathcal{D}_{z}}(a,b)=\mathbb{E}_{z\sim\mathcal{D}_{z}}\left[{O_{u}}(a,b,z)\right]\quad\text{where}\quad{O_{u}}(a,b,z)=\begin{cases}1/2\quad\text{if }u(a,z)=u(b,z)\\ \mathbf{1}\{u(a,z)>u(b,z)\}\quad\text{o.w.}\end{cases}$ (4)
@@ 2.2 Related Work
The challenges of preference learning that we group as cases of “hidden context” have also been studied individually. There has been some work on explicitly modeling annotator disagreement (Fleisig et al., 2023; Baumler et al., 2023) as well as other approaches to learning from annotators with diverse preferences (Jia et al., 2023; Dumoulin et al., 2023; Mishra, 2023; Fish et al., 2023). Other work has studied the effects of human irrationality or non-BTL models of human behavior on preference learning (Bobu et al., 2020; Lee et al., 2021; Laidlaw & Russell, 2021; Knox et al., 2022; Laidlaw & Dragan, 2022), which under our framework can be modeled as hidden context. Zhuang & Hadfield-Menell (2020) and Dai et al. (2023) study the optimization of multiple objectives learned from human preferences. Finally, related to our connections with social choice theory in Section 3, some previous work has associated preference or reward learning with concepts in economics, such as voting rules (Conitzer & Sandholm, 2005), incentive compatibility (Echenique & Prasad, 2019), and mechanism design (Fickinger et al., 2020).
@@ 3 Theoretical Analysis
BTL preference learning implicitly aggregates hidden context according to Borda count. That is, if $\hat{u}$ is optimized according to (3), then $\forall a,b\in\mathcal{A}$, $\hat{u}(a)>\hat{u}(b)\Leftrightarrow\text{BC}(a)>\text{BC}(b)$.
@@ 3.1 Comparison to expected utility and least-squares regression
The fact that least-squares utility regression converges to $\hat{u}=\bar{u}$ shows that, in some sense, it gracefully degrades in the presence of hidden context. Although there are problems with maximizing expected utility, it is a well-understood method of aggregating utilities over hidden contexts that has desirable decision-theoretic properties. Thus, it would be helpful if the utility function $\hat{u}(a)$ learned by preference learning with hidden context were equivalent to the expected utility $\bar{u}(a)$. In this section, we characterize when the output of preference learning with hidden context is equivalent to that of utility regression. We identify sufficient conditions for the two to be equivalent and present cases where they differ.
@@
Many noise distributions, such as uniform and normal distributions, clearly satisfy the assumptions of Theorem 3.2. Thus, as long as the noise caused by hidden context does not vary across alternatives and is not too unusual, we generally expect that preference learning will give a utility function with the same ordering over alternatives as the expected utility. This means that it performs similarly to least-squares regression.
@@
Let $F_{u\mid z}$ be the CDF of the distribution of $u(a,z)$ for some fixed value of $z$ when $a\sim\text{Unif}(\mathcal{A})$. Then for any $a$,
$\text{BC}(a)=\mathbb{E}_{z\sim\mathcal{D}_{z}}\left[F_{u\mid z}(u(a,z))\right].$
@@ 3.2 Connections to social choice theory
When training on comparison data from many agents, each with their own preferences, preference learning aggregates all their feedback into a single utility function. As we described in Section 2, this is a case where the identity of the annotator is hidden context: it affects the comparison outcomes, but is unseen by the preference learning algorithm. Social choice theory studies methods for aggregating preferences from a population. Thus, it can provide a lens through which to understand this particular case of preference learning with hidden contexts.
@@ 4 Distributional Preference Learning
Our alternative preference learning methods, which we call distributional preference learning (DPL), output a distribution over possible utilities for each alternative rather than a single value (Figure 3). In particular, we learn a mapping $\smash{\hat{\mathcal{D}}}:\mathcal{A}\to\Delta(\mathbb{R})$ from alternatives to distributions over utilities to estimate the distribution of $u(a,z)$ when $z\sim\mathcal{D}_{z}$. We consider two variants, each of which parameterizes the distribution $\smash{\hat{\mathcal{D}}}(a)$ in a different way.
First, the mean-and-variance model learns two functions $\hat{\mu}:\mathcal{A}\to\mathbb{R}$ and $\hat{\sigma}:\mathcal{A}\to[0,\infty)$, parameterizing the distribution over utilities as $\smash{\hat{\mathcal{D}}}(a)=\mathcal{N}\left(\hat{\mu}(a),\hat{\sigma}(a)^{2}\right)$. Second, in the categorical model, we choose $n$ evenly spaced utility values $u_{1}<u_{2}<\ldots<u_{n}$, and then parameterize the distribution as the probabilities of each of those utilities $\hat{p}(u_{i}\mid a)$ for $i\in\{1,\dots,n\}$. We train the distributional preference models by maximizing the likelihood of the data given the model $p_{\smash{\hat{\mathcal{D}}}}(a,b)=\mathbb{E}\left[{O}(u_{a},u_{b})\;\middle|\;u_{a}\sim\smash{\hat{\mathcal{D}}}(a),u_{b}\sim\smash{\hat{\mathcal{D}}}(b)\right]$. Concretely, for the mean-and-variance model, the loss for a single preference comparison where alternative $a$ is preferred to $b$ is the negative log probability that $u_{a}-u_{b}>0:$
@@ 5 Case Study: Competing Objectives in RLHF
In this section, we evaluate DPL’s ability to identify hidden context through a case study on large language model (LLM)-based reward models. Chatbots like GPT-4 and Claude are trained by learning a human reward model and then optimizing it via reinforcement learning, together referred to as RLHF. In order to evaluate the ability of DPL methods to identify hidden context, we use the HH-RLHF dataset (Bai et al., 2022a). For this dataset, raters were separately asked to provide preferences on whether responses were helpful or harmful. This allows us to determine if DPL can automatically detect this context.
@@ 6 Conclusion
Acknowledgments
We thank Ruiqi Zhong and Sam Toyer for feedback on drafts. Cassidy Laidlaw was supported by an Open Philanthropy AI Fellowship. Dylan Hadfield-Menell was supported by an AI2050 Early Career Fellowship from Schmidt Sciences.
@@ B.2 BTL and Borda Count
Given a set of preference $\{\succeq_{i}\}_{i=1}^{n}$, we call $\text{BC}:\mathcal{A}\to\mathbb{R}$ the Borda count:
$\text{BC}(a)=\sum_{i=1}^{n}\sum_{c\in\mathcal{A}}\textbf{1}\{a\succ_{i}c\}$